Calculate the area of a surface on which pressure exerted by a wooden block is 20kPa and weight of block is 100N


1
Expert's answer
2021-05-30T13:28:05-0400
"P=\dfrac{F}{A},""A=\dfrac{F}{P}=\dfrac{100\ N}{20\cdot10^3\ \dfrac{N}{m^2}}=0.005\ m^2."